The time it takes a mechanic to change the oil in a car is exponentially distributed with a mean of 5 minutes. What is the probability that it will take a mechanic up to 6 minutes to change the oil?
Answer #1

Let X be the time it takes for the mechanic to change the oil inthe car. Then X has an exponential distribution with parameterλ; X~Exp(λ=1/μ = 1/5 = 0.2).

P(X < 6) = 1 - P(X > 6) = 1 - e-λx = 1 -e-(0.2)(6) = 1 - 0.3011942 = 0.699